Hypothetical drug-development program requires $200 million in out-of pocket
costs over a 10-year period during which no revenues are generated, and with
only a 5% probability of success. However, if the drug development is successful, it
is plausible to assume that it could generate a net income of $2 billion per year over
a 10-year period of exclusivity from years 11–20. The present value of this income
stream in year 10 is $12.3 billion (using a 10% cost of capital).
(a) Compute the expected return and standard deviation (over a 10-year period) of
this investment
1
Expert's answer
2017-11-22T15:17:07-0500
E(R) = 0.05 * 61.5 + 0.95*0 = 3.075 (307.5%).
Return on capital (for successful project) = 12.3/0.2 = 61.5
Return on capital (for unsuccessful project) = 0/0.2 = 0
Standard deviation = (0.05*(12.3-0.615)2+0,95*(0-0.615)2 )1/2= 2.68.